OpenStudy (anonymous):

To solve the equation m plus five equals fourteen, you should: subtract 5 from both sides. subtract nineteen from both sides. add 5 to both sides. add fourteen to both sides.

OpenStudy (anonymous):

subtract 5 from each side m+5=14 m+5-5=14-5 m=9

OpenStudy (anonymous):

You should....Do your own homework maybe? The answer is Subtract 5 from both sides

OpenStudy (anonymous):

so its A

OpenStudy (anonymous):

ight thx mike

OpenStudy (anonymous):

Yep